OpenStudy (anonymous):

differentiate \[ y=x^{-1/2} \]

OpenStudy (anonymous):

please show me step by step

OpenStudy (anonymous):

i thought the answer would be -1/2...but its not right

OpenStudy (anonymous):

y=\[x ^{-1/2}\] dy/dx=-1/2*\[x ^{-1/2-1}\] and so u get -.5\[x ^{-3/2}\]

OpenStudy (anonymous):

differentiation of \[x ^{n}=n*x ^{n-1}\], so u can follow it up !

OpenStudy (anonymous):

mmm im trying to understand how youve got \[-1/2 x ^{-1/2-1)}\]

OpenStudy (anonymous):

the -1/2-1

OpenStudy (anonymous):

ok, see the general formula i gav. in the formula u can see that n=-1/2, so substitute u'll get answer, got it?

OpenStudy (anonymous):

y=x^-1/2 dy/dx= (-1/2)x^(-1/2 - 2/2) =(-1/2)x^(-3/2)

OpenStudy (anonymous):

ok..i get it thanks

OpenStudy (anonymous):

ok wc